# Entity Framing and Role Portrayal in the News

### v1

@@ Appendix A Detailed Taxonomy with Examples
Any references to “URW” and “CC” below denote the Ukraine-Russia War, and the Climate Change domains, respectively.
@@ A.2 Antagonist
Example: Figures involved in political scandals or espionage, such as Watergate conspirators or modern cyber espionage cases. In CC, this could manifest as persons or organizations conspiring to bypass environmental regulations to turn up a profit.
@@
Example: Politicians spreading false information for political gain, or media outlets engaging in propaganda.
@@ Appendix B Annotation Guidelines
We prepare these general set of guidelines to prepare the annotators and avoid human biases before starting the annotation:
  • The annotators should get acquainted with the two domains covered by the tasks; for instance,  Amanatullah et al. (2023) and Coan et al. (2021) provide a good coverage of the URW and CC domains,
@@
  • The annotators should not exploit any specific external knowledge bases for the purpose of annotating documents.
The annotation guidelines we prepare to annotate and curate the entity framing corpus are as follows:
Any references to “URW” and “CC” below denote the Ukraine-Russia War, and the Climate Change domains, respectively.
@@
    In this annotation task, the entities of interest are understood in a broad sense to include both traditional named entities (such as persons, organizations, and locations) as well as toponym-derived entities. Toponym-derived entities are phrases that indicate a group or collective identity based on a place or affiliation, including but not limited to:
@@
    The annotators are provided with a number of news articles and are expected to assign role(s) to named entities that are central to the article’s story according to the taxonomy of roles that was provided earlier.
@@
    The annotators are provided with a detailed taxonomy that includes definitions and examples.
@@
    For more details on what qualifies as a named entity, in addition to the definition of the broader sense of named entities in the first bullet point in these guidelines, the annotators should also examine the NER annotation guidelines outlined in http://www.universalner.org/guidelines/.
@@
    The annotators will pick one or more fine-grained roles for the named entities they believe are central to the article’s story.
@@
    Entity mentions can be assigned fine-grained roles from more than one main role. However, during curation, we will not be including these instances in the current version, even though we still annotate them.
@@
    The determination of how central a named entity is in an article is admittedly subjective. To reduce bias, such determination should be based on the careful reading of the article and the story it is pushing. An annotated example is provided in 1. Notice that named entities such as New York Times and Israel were not annotated because they are not central to the story.
@@
    As a general rule, annotators should annotate only the first mention per entity where it is clear that this entity has the specific role(s). There is no need to annotate subsequent mentions of this entity with the same role, But annotating more mentions with the same surface form and role is not a mistake, but it is simply not required.
    This rule also extends for surface mentions of the same entity. For example, “Putin” and “Vladimir Putin” are both surface mentions of the same entity, so only the first occurrence of any of the surface forms would be annotated.
    On the other hand, while entities such as “Moscow”, “Russia”, and “Putin” are closely related, they are not surface forms of the same entity, and are considered as distinct, separate entities.
@@
    Named entities appearing in indirect speech (e.g., quotes) should not be annotated. Indirect speech should be considered as supporting detail to the story, but named entities which are central to the narrative would likely appear also outside of quotations. This guideline helps avoid confusions inherent in indirect speech.
@@
    If the above would result in more than one mention of the same entity with the same role, the curator does not need to remove all these additional mentions. We keep all of them.
@@
    Should an entity mention that was previously annotated with a certain role appear in a different context with different role(s), the first mention where the role(s) changed should be annotated.
    The above rule is repeated for as many times as an entity changes roles across mentions. For example, if an en entity, let’s say NATO, appears 20 times in an article. The first 10 mentions show NATO as a Guardian and a Virtuous entity. The 11-15th mentions portray NATO as a Foreign Adversary, and the 16-20th mentions portray NATO as Exploited. Then we only need 3 annotations in total to account for the 3 different roles NATO was portrayed as. These 3 annotations should all be the first mention occurrences where NATO assumed each distinct set of roles (i.e., mention 1, mention 11, and mention 16 should be annotated).
@@
    Regarding scenarios where different surface forms for the same named entity (e.g., NATO vs North Atlantic Treaty Organization) appear in the article, it is sufficient to pick only one of the surface forms.
@@
    If the mention of the entity does not have any role vis-a-vis the taxonomy of roles then no role should be given. As a consequence, we do not need an “Other” label.
  16. 16.
    The curator may see conflicting annotations in the curation mode and could resolve the conflict, and then the remaining non-conflicting roles could be checked and adopted accordingly.
@@ Appendix C Experimental Settings
To prevent data leakage, we created train/dev/test splits based on entire articles rather than individual entity-mention annotations. The details of these splits are provided in Table 7.
BG EN HI PT RU All
Train 165 (389) 133 (440) 203 (1347) 206 (833) 89 (252) 796 (3261)
Dev 94 (237) 69 (245) 139 (983) 100 (417) 44 (114) 446 (1996)
Test 15 (30) 27 (90) 35 (279) 31 (115) 28 (85) 136 (599)
Total 274 (656) 229 (775) 377 (2609) 337 (1365) 161 (451) 1378 (5856)
@@ Appendix D Dataset Statistics
Figure 7: Top entities counts after multilingual linking was manually performed to link surface string to unique identifiers. The entities selected are all the ones for which at least one surface string has a count of a least 10 in any language
@@ Appendix F Annotation Tool
To annotate a mention of an entity with a role, annotators should go to the part of the article where the entity is mentioned and select it. After selecting an entity mention, annotators can then assign roles as shown in 12.

### v2

@@ Appendix A Detailed Taxonomy with Examples
Below, we definer the entries in our taxonomy, and we give an example for each category. Any references to “URW” and “CC” below denote the Ukraine-Russia War, and the Climate Change domains, respectively.
@@ A.2 Antagonist
Example: Figures involved in political scandals or espionage, such as Watergate conspirators or modern cyber espionage cases. In CC, this could manifest as persons or organizations conspiring to bypass environmental regulations for profit.
@@
Example: Politicians spreading false information for political gain, or media engaging in propaganda.
@@ Appendix B Annotation Guidelines
These guidelines aim to prepare the annotators and avoid human biases before starting the annotation:
  • The annotators should get acquainted with the two domains covered by the tasks; for instance,  Coan et al. (2021) and  Amanatullah et al. (2023) provide a good coverage of the CC and URW domains, respectively,
@@
  • The annotators should not exploit any external knowledge bases for the purpose of annotating documents.
Our guidelines for annotating and curating the entity framing corpus are as follows. Any references to “URW” and “CC” below denote the Ukraine-Russia War, and the Climate Change domains, respectively.
@@
    The entities of interest are understood in a broad sense to include both traditional named entities (such as persons, organizations, and locations) and toponym-derived entities. Toponym-derived entities are phrases that indicate a group or collective identity based on a place or affiliation, including, but not limited to:
@@
    Annotators are provided with a number of news articles and are expected to assign roles to named entities that are central to the article’s story, according to the taxonomy of roles that was provided earlier.
@@
    Annotators are provided with a detailed taxonomy that includes definitions and examples.
@@
    For more details on what qualifies as a named entity, in addition to the definition of the broader sense of named entities given above in these guidelines, the annotators should also examine the NER annotation guidelines in www.universalner.org/guidelines/.
@@
    Annotators pick one or more fine-grained roles for the named entities they believe are central to the article’s story.
@@
    Entity mentions can be assigned fine-grained roles from more than one main role. However, during curation, we will not include these instances in the current version of the corpus, even though we annotate them.
@@
    The determination of how central a named entity is in an article is admittedly subjective. To reduce bias, such determination should be based on the careful reading of the article.
@@
    As a general rule, annotators should annotate only the first mention of each entity where it is clear that this entity has the specific role(s). There is no need to annotate subsequent mentions of this entity with the same role, but annotating more mentions with the same surface form and role is not a mistake; it is simply not required.
    This rule also extends to surface mentions of the same entity. For example, “Putin” and “Vladimir Putin” are both surface mentions of the same entity, so only the first occurrence would be annotated.
    On the other hand, while entities such as “Moscow”, “Russia”, and “Putin” are closely related, they are not surface forms of the same entity, and are considered to be distinct separate entities.
@@
    If the above results in more than one mention of the same entity with the same role, the curator does not need to remove all of these additional mentions. We keep all of them.
@@
    Should an entity that was previously annotated with a certain role appear in a different context with different roles, the first mention where the roles changed should be annotated.
    The above rule is repeated as many times as the entity changes roles across mentions. For example, if an entity, let’s say NATO, appears 20 times in an article, the first 10 mentions show NATO as a Guardian and a Virtuous entity. The 11th-15th mentions portray NATO as a Foreign Adversary, and the 16th-20th mentions portray NATO as Exploited, then we need only 3 annotations in total to account for the 3 different roles that NATO was portrayed as. These 3 annotations should all be the first mentions where NATO assumed each distinct role (i.e., mention 1, mention 11, and mention 16 should be annotated).
@@
    If different surface forms for the same named entity (e.g., NATO vs. North Atlantic Treaty Organization) appear in the article, it is sufficient to annotate only one of the surface forms.
@@
    There is no “Other” label in the taxonomy, as mentions without a discernible role in relation to the taxonomy are simply not assigned any role.
@@
    The curator may see conflicting annotations in the curation mode and can resolve the conflict, and then the remaining non-conflicting roles can be checked and adapted accordingly.
@@ Appendix D Dataset Statistics
Figure 7: The 10 most frequent (a) main role and (b) fine-grained role transition sequences.
@@ Appendix F Annotation Tool
To annotate a mention of an entity with a role, annotators should go to the part of the article where the entity is mentioned and select it. After selecting an entity mention, annotators can then assign roles as shown in 13.